OpenStudy (anonymous):

How much material will be needed for the rungs of a ladder of 31 rungs, if the rungs taper uniformly from 18 in to 28 in? Identify the nature of the sequence and present algebraic analysis to answer the question asked.

OpenStudy (cwrw238):

first term of sequence = 18 in and the next 30 terms (rungs) taper uniformly up to 28 in this is an arithmetic sequence sum of n terms = (n/2)[ a + l) where a = first term and l = last term so for 31 terms we get total material needed = (31/2)(18 + 28) ins

OpenStudy (anonymous):

thank you! clearly.